Solve for x
Y=x^2-6x+8
2y+x=4
Solve for x
X^2+1-y=0
X-1-y=0

Answer by stanbon(75887)

Solve for x
Y=x^2-6x+8
2y+x=4
------
x = 4-2y
Substitute for "x" and solve for "y":
y = (4-2y)^2-6(4-2y)+8
y = 16-16y+4y^2 -24 + 12y + 8
4y^2 -5y = 0
y(4y-5) = 0
y = 0 or y = 5/4
solve for "x":
If y = 0, x = 4-2*0 = 4
If y = 5/4, x = 4-2(5/4) = 4-(5/2) = 3/2
================================================
Solve for x
X^2+1-y=0
X-1-y=0
=============
y = x-1
Substitute for "y" and solve for "x"::
x^2 -(x-1) + 1 = 0
x^2 -x +2 = 0
Factor:
x = [1 +- sqrt(1 - 4*2]/2
---
x = (1/2) +- isqrt(7)/2
-----
Since y = x-1, y = (-1/2) +- isqrt(7)/2
============================================
Cheers,
Stan H.
==========================